### v3

# Power sector effects of alternative options for de-fossilizing heavy-duty vehicles: go electric, and charge smartly

Abstract
In the passenger car segment, battery-electric vehicles (BEV) have emerged as the most promising option to de-fossilize transportation. For heavy-duty vehicles (HDV), the technology space still appears to be more open. Aside from BEV, electric road systems (ERS) for dynamic power transfer are discussed, as well as indirect electrification with trucks that use hydrogen fuel cells or e-fuels. Here we investigate the power sector implications of these alternative options. We apply an open-source capacity expansion model to future scenarios of Germany with high renewable energy shares, drawing on detailed route-based truck traffic data. Results show that power sector costs are lowest for flexibly charged BEV that also carry out vehicle-to-grid operations, and highest for HDV using e-fuels. If BEV and ERS-BEV are not charged in an optimized way, power sector costs increase, but are still substantially lower than in scenarios with hydrogen or e-fuels. This is a consequence of the poor energy efficiency of indirect HDV electrification, which outweighs its temporal flexibility benefits. We further find a higher use of solar photovoltaic energy for BEV and ERS-BEV, while hydrogen and e-fuel HDV lead to a higher use of wind power and fossil electricity generation. Results are qualitatively robust in sensitivity analyses without the European interconnection or with capacity limits for wind power expansion.
keywords:
Heavy-duty vehicles, battery-electric vehicles, catenary, hydrogen, power sector modeling
[…]
[DIW]organization=German Institute for Economic Research (DIW Berlin), addressline=Mohrenstraße 58, city=Berlin, postcode=10117, country=Germany
[…]
[ifeu]organization=ifeu - Institute for Energy and Environmental Research, addressline=Wilckensstraße 3, city=Heidelberg, postcode=69120, country=Germany
1 Introduction
Making energy consumption climate neutral in all end-use sectors is of paramount importance for mitigating climate change de_coninck_strengthening_2018. A key strategy for achieving this is to substitute fossil fuels by renewable electricity, facilitated by direct or indirect electrification of end uses in mobility, heating, and industrial applications shukla_contribution_2022. In the transportation sector, battery-electric vehicles (BEV) have emerged as the most promising option for de-fossilizing the passenger car segment. Already today, BEV can lead to sizeable greenhouse gas emission reductions compared to internal combustion engines hoekstra_underestimated_2019, which will further increase when the electricity mix becomes cleaner. In many countries, markets for electric passenger cars have been soaring in the past years, and are expected to continue to grow strongly in the near future iea_global_2022. For heavy-duty vehicles (HDV), however, the technology space still appears to be more open. While the feasibility of pure battery-electric HDV has been assessed to be increasing nykvist_feasibility_2021, they compete with other options. This includes electric road systems (ERS), which allow for dynamic power transfer to electric vehicles on the road (boltze_insights_2020; speth_comparing_2021); trucks with hydrogen fuel cells; or conventional HDV with internal combustion engines that use power-to-liquid (PtL), also referred to as e-fuels, which are produced mostly with renewable electricity hannula_nearterm_2019; lajevardi_simulating_2022; plotz_hydrogen_2022; li_transition_2022.
These options of direct or indirect electrification of HDV have different properties concerning, on the one hand, energy efficiency, and, on the other hand, temporal flexibility of electricity use. For example, direct electrification via BEV is more energy efficient compared to indirect electrification via electrolysis-based hydrogen or e-fuels ueckerdt_2021; lajevardi_simulating_2022. Yet, the temporal flexibility of BEV may be constrained by charging availability and limited battery capacities, as vehicle batteries are costly and heavy. In contrast, indirect electrification via hydrogen or e-fuels may entail large-scale and low-cost storage options taljegard_2017; stoeckl_2021, but the overall energy efficiency of these supply chains is lower compared to BEV. Temporal power sector flexibility becomes increasingly important with growing shares of renewables, as the potential for firm renewable generation such as hydropower, bioenergy, or geothermal power is limited in many countries. In contrast, wind and solar power potentials are often abundant, but they have variable generation profiles that depend on weather conditions and daily and seasonal cycles (lopezprol_economics_2021). Integrating growing shares of variable renewable energy sources thus requires an increasing use of temporal and spatial flexibility options in the power sector. Flexibility may be provided on the supply side (e.g., dispatchable generators) or the demand side (e.g., flexible loads), and via energy storage or transmission (kondziella_flexibility_2016; schill_joule_2020).
Against this background, we investigate the power sector implications of different options for (in-)directly electrifying HDV, particularly focusing on the trade-off between energy efficiency and temporal flexibility. To do so, we apply an open-source capacity expansion model zerrahn_long-run_2017; gaete_2021 to scenarios of the Central European power sector for the year 2030 with high renewable energy shares. We focus on the domestic traffic of HDV in Germany with a gross vehicle weight above 26 tonnes, drawing on a detailed data set of truck trips on inner-German routes. We include stationary-charged BEV trucks as well as hybrid battery-catenary trucks as a particular example of an electric road system technology (ERS-BEV) [^1] In general, there are three options available for ERS: Overhead catenary lines, conductor rails, and inductive power transfer. We have opted for the first option as it is the technology currently under discussion in Germany with three test tracks in place (Hacker2023)., fuel-cell hydrogen electric trucks (FCEV), and such with internal diesel combustion engines powered by e-fuels (ICEV PtL), see Figure 1. For hydrogen and e-fuels, we assume domestic supply chains. We compare the power sector costs [^2] Power sector costs comprise fixed and variable costs of all electricity generation and storage technologies over a full year. They do not include the costs of HDV electrification infrastructure, e.g., overhead lines, charging stations, or filling stations for hydrogen or PtL. of these options, as well as their repercussions on the optimal power plant fleet and its use.
Figure 1: Overview of direct and indirect HDV electrification options covered in this analysis. For direct electrification via BEV and ERS-BEV we differentiate scenarios with flexible charging without and with V2G, as well as scenarios with inflexible, i.e., non-optimized charging. Hydrogen and PtL options are assumed to be always operated flexibly, constrained by hydrogen or e-fuel storage capacities, and do not allow reconversion to electricity.
While there is a broad literature on the potential power sector impacts of battery-electric passenger cars richardson_electric_2013; muratori_shape_2020; mangipinto_impact_2022, according research for electric HDV is sparse. A common finding of analyses focusing on passenger cars is that controlled charging can help to avoid problematic load peaks, to integrate renewable electricity generation, and to lower overall system costs schill_power_2015; gnann_2018; sadeghian_a_2022. In particular, feeding back electricity from car batteries to the system (vehicle-to-grid, V2G) can help to balance the diurnal variability of solar photovoltaic (PV) generation brown_2018.
How electrified HDV interact with the power sector is hardly explored so far. Some analyses investigate grid or peak load impacts without using power sector models. For example, an analysis of the distribution grid impacts of BEV-HDV with depot charging finds that many existing substations in the US would allow for massive HDV charging without grid upgrades borlaug_heavyduty_2021. Earlier studies conclude that local grid impacts of electrified HDV near logistics centres and along major roads may be substantial in Switzerland and Finland (liimatainen_2019), or that ERS-HDV could strongly increase the peak load in Norway taljegard_2017. Using capacity expansion power sector models, other analyses come to deviating conclusions on the impacts of electric HDV on the overall load profile and the wider power sector; in a Scandinavian-German case study that includes ERS-HDV in a subset of scenarios, effects are pronounced taljegard_impacts_2019, while in a European model study the impacts are moderate plotz_impact_2019. Importantly, both analyses assume exogenous, i.e., inflexible, charging profiles of HDV and do not temporally optimize their (dis-)charging operations. Given this research gap, we deliberately focus our analysis on the power sector effects of electrified HDV, modelling in detail their repercussions on optimal capacities and their hourly use. In doing so, we complement other strand of the literature which have focused on the total costs of ownership (TCO) nykvist_feasibility_2021; hunter_spatial_2021; samet_2024, energy efficiencynoll_analyzing_2022, or potential health and climate impacts of electric HDV lin_2021.
We contribute to the literature with a dedicated analysis of the power sector impacts of different HDV electrification options that co-optimizes their charging and discharging operations (including V2G) with the capacity and dispatch decisions in the power sector. We do so for a wide range of HDV technologies, including dynamic power supply via electric road systems. We use a power sector model that fully captures the hourly variability of load and renewable generation over all hours of a full year, and apply it to a future scenario with high shares of variable renewables. The model code and all input data, including detailed hourly HDV mission profiles for domestic transport in Germany, are provided open-source for transparency and reproducibility (see section 4).
2 Results
In the main part of this paper, we show results for a setting where Germany is interconnected with its neighbors, defined as the baseline model specification. Each country is modeled as a single node. While this configuration can be considered to be a policy-relevant scenario for 2030, where electric HDV in Germany may benefit from the flexibility provided by the European interconnection, we also provide results for a sensitivity analysis where the German power sector is modeled in isolation in the Supplemental Information (SI.4.1). [^3] In the SI, we also show further results for baseline assumptions (section SI.3), a sensitivity analysis that explores the effects of an upper bound for wind power capacity (section SI.4.2), and another sensitivity with alternative charging availability assumptions (section SI.4.3).
We compare nine HDV scenarios against a reference without any electrification of HDV (Ref). Three scenarios cover pure BEV trucks with varying degrees of flexibility: these are either charged as flexibly as possible (BEV Flex), additionally have the option of feeding back electricity to the grid (BEV Flex V2G), or are charged in a way that is not co-optimized with the power sector (BEV Inflex). There are three respective scenarios for ERS-BEV, which can also draw electricity from overhead lines and have smaller batteries than pure BEV. Two scenarios cover fuel-cell trucks with either centralized electrolysis and transport via liquefied hydrogen, which enables low-cost hydrogen storage (FCEV Centralized), or temporally inflexible on-site electrolysis at filling stations (FCEV On-Site). The last scenario represents conventional trucks with internal combustion engines powered by e-fuels (ICEV PtL). To separate effects, we assume that the whole domestic fleet of HDV (318,700 vehicles > 26 t) in the year 2030 consists of the particular vehicle type (compare 1). While this is not a policy-relevant assumption for 2030, our aim is to illustrate corner solutions. The power sector effects of mixed real-world HDV fleets are likely between the ones of the extreme cases considered here.
2.1 Lowest power sector costs and electricity prices for BEV with V2G
Figure 2: Changes in yearly power sector costs and electricity demand induced by different HDV options (upper panel), and average wholesale market prices of charging electricity (lower panel).
Compared to the reference case without electrified trucks, yearly power sector costs increase in all scenarios with electrified HDV (Figure 2, upper panel). The cost of the additional electricity demand induced by HDV thus always outweighs their potential flexibility benefits. Cost effects, however, vary considerably between different options. Flexible BEV with V2G incur the lowest additional power sector costs of 1.8 bn Euros/year (around 5,600 Euros/year per vehicle), followed by optimally charged BEV without V2G with 2.3 bn Euros/year (around 7,200 Euros/year per vehicle). If BEV charging is not optimized, system costs are markedly higher, with 3.8 bn Euros/year (11,900 Euros/year per vehicle). Results are qualitatively similar for ERS-BEV, but on a slightly higher cost level. The differences between the three ERS-BEV cases are less pronounced than for pure BEV, as their temporal flexibility potential is much smaller. The battery capacity of an ERS-BEV is only around a quarter of that of a pure BEV (655 kWh usable capacity per truck vs. 181 kWh), so ERS-BEV can make less use of hours with low electricity prices (compare also Figure SI.9). In contrast, power sector costs are higher for FCEV (12.6 or 12.7 bn Euros/year, for decentralized or centralized hydrogen provision, i.e. around 39,700 Euros/year per vehicle) and even more so for PtL (16.8 bn Euros/year, or 52,700 Euros/year per vehicle). This is a consequence of high conversion losses of hydrogen and PtL supply chains and vehicles drive trains. Because of these losses, the two hydrogen supply chains increase the electricity demand more than twice as much as the battery-electric options. The electricity demand of PtL-HDV is nearly four times as high as in the case of BEV. Notably, the cost differences between BEV and ERS-BEV scenarios are much smaller than the differences between these direct-electric options and indirect electrification via hydrogen or PtL.
We also evaluate average yearly wholesale electricity prices for HDV electricity (Figure 2, lower panel). [^4] We interpret the dual variables of the model’s energy balance as hourly wholesale prices, which is common practice in energy modeling (brown_decreasing_2021; lopezprol_economics_2021). This allows to separate the differences in overall electricity consumption of the various HDV options from their ability to make use of low-cost electricity. Average prices are calculated by multiplying hourly wholesale prices of electricity consumed by the different HDV options or fed back to the grid with respective hourly quantities, summing up over the whole year, and dividing by the overall electricity consumption of the fleet. That is, the numbers also account for revenues of electricity sold via V2G, assuming that HDV operators receive the respective hourly wholesale price whenever they feed electricity back to the grid. BEV with V2G accordingly face the lowest average electricity prices, as these also benefit from revenues of feeding back to the grid, followed by BEV without V2G.
Average electricity prices paid by ERS-BEV are slightly higher, as their smaller batteries limit the ability to optimize their charging and V2G decisions. In contrast, pure BEV can leverage their larger batteries to make better use of hours with low electricity prices. In the case of inflexible charging, average electricity prices faced by inflexibly charged BEV are high, and even slightly above those of inflexible ERS-BEV. This is because inflexibly charged BEV benefit less from cheap electricity prices around midday related to abundant PV feed-in than inflexible ERS-BEV, which on average have a better (catenary) grid connection during these hours.
Prices for electricity used in hydrogen and PtL supply chains are in the same range as those of ERS-BEV options but below those of inflexibly charged BEV or ERS-BEV. Centralized hydrogen and PtL supply can make use of lower prices than decentralized supply, because their low-cost storage options allow for higher temporal flexibility. Electricity prices of centralized hydrogen supply chains and PtL are also cheaper than those faced by inflexibly operated BEV or ERS-BEV. In terms of overall costs, these temporal flexibility benefits are, however, by far outweighed by the higher overall energy consumption of the FCEV and PtL options (Figure 2, upper panel).
The average wholesale price in the reference scenario without electrified HDV is 84.9 Euros/MWh. Inflexibly charged BEV and ERS-BEV, as well as inflexible on-site electrolyzers, face slightly above-average prices. In contrast, the more flexible HDV options can make use of below-average market prices.
2.2 Capacity and dispatch effects
Figure 3: Effects of different HDV scenarios on optimal generation capacity (upper panel) and on yearly generation (lower panel) in Germany.
The upper panel of Figure 3 shows optimal generation capacities in the reference (on the left) and the changes induced by HDV (on the right) in Germany. In the reference, variable solar PV and onshore wind power dominate the capacity mix. [^5] In the Supplemental Information, we discuss the effects of including an upper bound for wind power investments by the year 2030, see section SI.4. These are complemented by smaller firm capacities of natural gas and bioenergy. Optimal capacity additions related to the electrification of HDV are predominantly a mix of solar PV and onshore wind power, as these are the lowest-cost options available. Flexible BEV and ERS-BEV lead to the highest PV shares in the capacity additions, especially if combined with V2G. BEV trucks with V2G serve as a short-duration electricity storage option with a low energy-to-power ratio and many yearly cycles, which favors the expansion of solar PV. The potential for solar PV integration via V2G is less pronounced for ERS-BEV than for BEV, because ERS-BEV have smaller batteries.
[…]
Yearly electricity generation in the reference, as well as the changes induced by different HDV scenarios, are shown in the lower panel of Figure 3. The picture is generally similar to optimal capacity changes, but the share of wind power in additional electricity generation tends to be higher because of higher full-load hours as compared to solar PV. Flexible BEV, particularly if combined with V2G, cause the highest increase in PV generation, which can increase the renewable share of the German power sector to nearly 84% (section SI.3.3). Centralized hydrogen or PtL supply chains, which require much more electricity than BEV and ERS-BEV, use their temporal flexibility to also increase electricity imports and the use of the most energy-efficient natural gas-fired power plants (CCGT). Note that this flexibility comes from the option of low-cost hydrogen or PtL storage. This is different in the case of temporally inflexible hydrogen production at filling stations (FCEV On-Site). As imports and CCGT are less available in the hours of decentralized hydrogen production, more wind and solar generation in combination with long-duration electricity storage are used, which also increases the share of renewables in the overall system to nearly 83%. Higher utilization of natural gas also translates into increasing direct carbon emissions (section SI.3.4). The effects of BEV and ERS-BEV on CO2 emissions are generally lower, and may even become negative when combined with V2G. This is because V2G allows for an increase in solar photovoltaic capacity and a higher renewable energy share, which can slightly reduce overall carbon emissions compared to the reference case without electrified HDV.
2.3 Time series reveal differences in flexibility characteristics
Figure 4: Electricity time series sample of five exemplary days (unit: GW): “HDV electricity demand” illustrates electricity flowing to BEV or ERS-BEV batteries, including amounts later charged back via V2G, or electricity needs of hydrogen and PtL supply chains; “HDV discharging to the grid” are V2G power flows from BEV or ERS-BEV to the grid; and “Residual load” is the net load of the electricity sector after considering the feed-in potential of variable renewables. The samples start on a Saturday. For simplification, we assume that both Saturday and Sunday are completely truck-free. BEV Flex and ERS-BEV Flex without V2G are omitted in the graph, as charging patterns are very similar to the shown V2G cases. Additional time series are shown in the SI.
Flexibly operated BEV trucks (BEV Flex V2G) are able to charge their batteries in hours of low residual load, as illustrated in Figure 4. Such periods generally go along with low wholesale market prices (compare Figure SI.9). This enables BEV to utilize renewable surplus energy (i.e., negative residual load) to a substantial extent. They also make use of the V2G option and feed back renewable surplus energy to the grid, whenever the battery capacity is not needed for driving. This is particularly visible during summer during weekends, when trucks are not driving, but assumed to have a grid connection (top right panel in Figure 4). Here, BEV with V2G charge their batteries with solar PV surplus energy on Saturday afternoon, and feed it back to the grid in the night between Saturday and Sunday. This is different in the following days, as HDV are used between Monday and Friday, and much less battery capacity is available for V2G. If BEV charging is not optimized, but follows a pre-determined pattern (BEV Inflex), charging profiles are less peaky and much more balanced. BEV are not able to make particular use of cheap renewable surplus electricity in this case, but carry out a substantial part of their charging in hours with positive residual load.
ERS-BEV follow similar patterns as non-catenary BEV. Yet, their smaller batteries make them temporally less flexible, so they can make less use of renewable surplus energy and have to draw more electricity from the grid during hours with positive residual load. Their potential for feeding electricity back to the grid is thus also much smaller than in the case of pure BEV. ERS-BEV sometimes also charge their batteries during driving (compare Figure SI.10).
Hydrogen and PtL supply chains show very different patterns of electricity use. Centralized hydrogen supply chains (FCEV Centralized), which come with large-scale hydrogen storage capacities, generally have a flat consumption profile in many hours, as their high fixed costs make it optimal to use them with high full load hours. This limits their ability to make use of renewable surplus energy. Yet, they can use the temporal flexibility provided by large-scale hydrogen storage to reduce electricity consumption in hours of high positive residual load, i.e., high prices. Decentralized hydrogen supply chains (FCEV On-Site) have an electricity use profile that largely follows the hydrogen demand profile, driven by limited on-site hydrogen storage capacity. Accordingly, their potential for avoiding electricity consumption in hours of high residual load is lower. The PtL supply chain (ICEV PtL) also comes with large-scale storage and thus has a similar pattern as the centralized hydrogen option. The load, however, is generally higher because of higher overall energy consumption of PtL versus hydrogen. This goes along with larger investments in electrolyzers, which are 24.4 GW in the ICEV PtL scenario, as compared to 13.5 GW for FCEV Centralized (section SI.3.2).
3 Discussion
3.1 Limitations and avenues for future research
Our model-based analysis of the power sector effects of alternative options for de-fossilizing heavy-duty vehicles inevitably has a range of limitations. We briefly discuss how these may qualitatively impact our results.
First, we deliberately focus on a detailed analysis of the power sector implications of alternative HDV electrification options to address this particular research gap. Unlike other studies, we do neither aim to compare total costs of ownership (nykvist_feasibility_2021; hunter_spatial_2021; samet_2024), nor the overall system costs of different options. Including the latter in our power sector-focused analysis would also require considering infrastructure costs related to catenary lines, chargers, or filling stations, as well as differences in vehicle costs. Such future costs are highly uncertain, so Monte Carlo approaches or other methods to deal with parameter uncertainty would have to be used. Note that the differences in the costs of vehicles and related charging infrastructure might in fact partially outweigh the differences in power sector costs identified here, especially the small differences between pure BEV and ERS-BEV. Yet, the very large power sector cost differences between (ERS-)BEV on the one hand, and FCEV or PtL trucks on the other, would require the latter to be substantially cheaper in terms of vehicle and infrastructure costs in order to break even, which appears unlikely.
While focusing on the power sector, we further assume extensive charging and overhead line infrastructure to be available for BEV and ERS-BEV, both in depots and en route, and that vehicles are always connected to the grid when idling. In reality, the availability of grid connections is likely to be lower for a variety of reasons, which would decrease the flexibility potentials of optimized charging. In sensitivity analyses, we show that decreasing charging availability at depots reduces the potential for flexible charging and particularly for vehicle-to-grid operations of BEV HDV (SI.4.3). This appears to be in line with the finding that the flexibility benefits of very high levels of grid connection may not necessarily outweigh related infrastructure costs (brinkel_2020). In further sensitivities, we conversely find that higher assumptions on fast-charging power ratings do not have a major effect on results.
Another simplification of our numerical analysis is the assumption of perfect foresight, which results in an upper limit of the flexibility benefits that can be realized by optimized (dis-)charging of BEV and ERS-BEV. We also abstract from including electric passenger vehicles or other future sector coupling options, which may provide flexibility to the power sector and compete with electric trucks for electricity use in times of high renewable availability. In particular, power-to-heat technologies may be able to efficiently exploit hours of high renewable generation and low energy prices, as heat storage is relatively cheap compared to electricity storage (bloess_2018). In contrast, we may underestimate temporal flexibility constraints of the PtL production process.
To facilitate numerically efficient solutions and keep results tractable, we further assume electricity generation capacities to be fixed in Germany’s neighboring countries. At the same time, the modeled HDV fleets only cover domestic truck transport in Germany. This may lead to an overestimation of the power sector flexibility supplied by Germany’s neighbors for integrating electric trucks in Germany. Likewise, increasing the spatial resolution within a country could raise additional insights. Taking into account transmission grid bottlenecks and differences in regional potentials for wind and solar power, it may be more beneficial to charge HDV trucks in some regions than in others. We further abstract from comparing different options for the distribution of hydrogen to filling stations, which has been done in detail in previous work (stoeckl_2021).
Future research may address these limitations. While our analysis focuses on the power sector, it also appears desirable to investigate complementary cost measures. This includes detailed total cost of ownership analyses from a HDV operator perspective, as well as investigations of overall system cost effects, which would also consider cost differences of charging and ERS infrastructures, as well as purchase cost differences of different types of HDV. This requires a whole set of additional and detailed parameter assumptions, which we leave for future research. Further, it would be desirable to carry out similar analyses for other countries and world regions. While we focus on Germany, we expect that general results should also hold for comparable non-hydro electricity sectors in other countries in temperate climates.
3.2 Differences in temporal flexibility and energy efficiency drive effects
Our analysis focuses on the effects of alternative options for de-fossilizing heavy-duty vehicles on power sector costs as well as optimal investment and dispatch decisions of electricity generation and storage technologies. We find that differences in temporal flexibility and energy efficiency are important drivers of results.
Fleets of BEV or ERS-BEV trucks that are charged in an optimized way cause the lowest increase in power sector costs. This is because they require much less electricity than HDV powered by electrolysis-based hydrogen or e-fuels. At the same time, they can offer temporal flexibility to the power sector, which lowers the costs compared to non-optimized charging. This is increasingly valuable with growing shares of variable renewable energy sources. Flexibility benefits are particularly large if BEV also carry out vehicle-to-grid operations, i.e., if the truck batteries can be used as a distributed grid storage option. In contrast, indirect electrification of the HDV fleet via FCEV or PtL incurs much higher power sector costs, even if compared to inflexibly charged BEV or ERS-BEV. That is, the temporal flexibility benefits of centralized hydrogen or PtL supply chains, which come with low-cost storage options, do not outweigh their energy efficiency drawbacks in terms of power sector costs.
[…]
Additional sensitivity analyses show that results are generally robust. For instance, Germany is modeled as an electric island in a sensitivity analysis, i.e., the flexibility benefits of the European interconnection are neglected (section SI.4.1). In such a case, flexible BEV with V2G are even more beneficial, and inflexible BEV operations lead to slightly higher cost increases because an isolated power sector is more flexibility-constrained. Further, results are robust against the assumption that the potential for wind power expansion is limited, and solar PV accordingly plays a larger role (section SI.4.2).
3.3 Direct electrification and smart charging preferable
From a pure power sector perspective, direct electrification of the truck fleet would thus be clearly preferable. Aside from the cost benefits determined here, this appears to be even more important as the utilization of renewable energy sources is unlikely to grow exponentially (hansen_2017), and renewable growth rates in fact have been smaller than what would be required in $1.5^{\circ}C$-compatible scenarios cherp_2021. A similar argument has been made for the global scale-up of electrolysis (odenweller_2022). All of this calls for energy-efficient, direct electrification options of HDV wherever possible.
Moreover, smart charging of BEV or ERS-BEV is highly desirable, as it provides temporal flexibility which leads to lower power sector costs and CO2 emissions, compared to non-optimized charging. This corroborates the findings of Pickering et al., who also highlight the benefits of controlled vehicle charging in decarbonized and largely renewable future European energy systems pickering_diversity_2022. If feeding back electricity to the grid via V2G is possible, this further lowers the costs and increases the optimal shares of solar PV.
4 Experimental procedures
4.1 The power sector model DIETER
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER). It is a linear program that minimizes power sector costs by optimizing capacity and dispatch decisions for a full year in an hourly resolution zerrahn_long-run_2017; gaete_2021. Its objective function includes fixed an variable costs of all electricity generation and storage technologies, electrolysis and PtL plants, as well as hydrogen or e-fuels transportation. It does not include the costs of charging or catenary infrastructure, hydrogen filling stations, or PtL filling stations. Accordingly, the power sector cost figures provided above do not include the costs of HDV electrification infrastructure. We further do not consider the option of hydrogen imports, as these are likely to be unavailable at scale by 2030. In general, the global scaling up of green hydrogen supply remains uncertain odenweller_2022.
Endogenous model variables include power sector costs, optimal generation and electricity storage capacities (Germany) and their hourly use (all countries), hourly decisions for HDV charging and discharging, as well as the capacity and operational decisions of electrolysis and PtL generation and storage infrastructure. In addition, we interpret the marginals of the hourly energy balance as wholesale prices (brown_decreasing_2021).
Exogenous model inputs include fixed and variable costs of all electricity generation and storage technologies, efficiency parameters, as well as time-series of variable renewable energy availability and electric load. We further include HDV driving profiles as explained in the following section 4.2 and section SI.1. In the case of inflexible HDV charging (BEV Inflex, ERS-BEV Inflex), we assume that the vehicles always start charging as soon as an opportunity arises, and that vehicles batteries are fully charged by the time the next trip starts, if possible. That is, the charging power is lower, the longer a vehicle is connected to the grid. This resembles the “balanced” charging profile defined in gaete_2021_emobpy.
The geographic scope of the model version used here includes Germany and its neighboring countries plus Italy. In order to reduce numerical complexity and improve tractability, we allow for endogenous generation capacity investment only in Germany, and fix the power plant portfolio for the other countries to values derived from ENTSO-E’s Ten Year Network Development Plans tyndp_2018; tyndp_2020. We further assume upper limits for investments into fossil generation technologies according to the main scenario for 2030 from the German Federal Grid Development plan (NEP, nep_2018).
Different versions of the model have been used in various earlier studies to investigate various aspects of renewable energy integration, electricity storage, and sector coupling (e.g.,  schill_2018; schill_joule_2020; stoeckl_2021; gils_2022). The DIETER model is available open-source. [^7] General DIETER repo: https://gitlab.com/diw-evu/dieter_public. The particular model version used for the present study, together with the complete input data set (including BEV and ERS-BEV profiles), is available under permissive licenses in a dedicated repository. This also includes model results in Jupyter notebook and html formats: https://gitlab.com/diw-evu/projects/my_eroads. A general model documentation is available here: https://diw-evu.gitlab.io/dieter_public/dieterpy/.
4.2 Traffic data of heavy-duty vehicles
We generate synthetic truck usage patterns that are intended to approximate the German fleet of HDV larger than 26 tons. The main data source for the usage patterns is the traffic model PTV Validate (ptv_2023), from which we extract a database of daily truck trips in domestic German road freight transport. In order to derive typical daily driving profiles, the following steps are performed (for more detailed information, see SI):
  1. 1.
    Calculation of annual mileage for transport relations. For this step, we build on statistics of typical annual mileages for different distance classes and derive a steady function.
  2. 2.
    Calculation of daily idle time. This is calculated based on the number of daily truck trips (and the number of stops in between) as well as typical times for loading and unloading, depending on the goods type. In addition, a mandatory driver’s break of 45 minutes is assumed for daily driving profiles > 4.5 hours.
  3. 3.
    The daily operating time of a vehicle on a particular route is determined from the above assumptions using the formula “operating time = journey time + idle time + driver’s break”.
Repeating steps 1-3, synthetic daily travel profiles with a certain daily operating time are obtained for all transport relations in the model. In order to limit these to a manageable number of daily driving profiles for the power sector model, all relations that have approximately the same daily operating time are combined. We derive a number of 19 daily profiles which are distributed over the course of the day such that the empirically observed course of the daily mileage on the German road network is reproduced. For these profiles, the time series of charging availability (in the depot, during idle and driver’s resting times) and of the electricity demand of pure BEV-HDV (500 km battery range) and ERS-HDV (150 km battery range) are calculated. The resulting electricity demands and charging availabilities are used as inputs for the DIETER model.
For FCEV and PtL trucks, only aggregate fuel demand time series are required. These are based on the same empirical traffic data from German highways (bast_2023). For FCEV, we assume a specific hydrogen demand of 6.8 kg/100km, and for ICEV-PtL a specific Diesel demand of 27.1 l/100km.
Acknowledgments
We thank Hinrich Helms and Michel Allekotte for their valuable comments on earlier drafts. This work has benefited from research grants by the Federal Ministry for the Environment, Nature Conservation and Nuclear Safety (BMU) and by the Federal Ministry for Economic Affairs and Climate Action via the projects “My eRoads” (Fkz 16EM4006-1) and “enERSyn” (Fkz 01MV22004B).
Author contributions
Conceptualization: JJ, WPS; Methodology: CG, FH, JJ, WPS; Software: CG; Investigation: CG, FH, JJ, WPS; Data curation: FH, WPS; Writing - original draft: CG, WPS; writing - review & editing FH, JJ; Visualization: CG, FH; Project administration: JJ; Funding acquisition: JJ, WPS.
Resource availability
The DIETER model version used for the present study, including input data as well as model results, are available under permissive licenses in a public repository: https://gitlab.com/diw-evu/projects/my_eroads.